OpenStudy (anonymous):

Find the exact area under the graph f(x)=x^2+5 from x=0 to x=2 by taking the limit of the Rieman Sum using right endpoints.

OpenStudy (abb0t):

So your bounds are [b,a] and you were given [0, 2] find your ∆x. Which is \[\frac{ b-a }{ n }\]

OpenStudy (anonymous):

Ok so the delta x is 1.

OpenStudy (abb0t):

Plug it into your equation.

OpenStudy (anonymous):

Wait should i use 2 as my n value since it doesnt specify how many rectangles

OpenStudy (abb0t):

use n, because you aren't given a specified vaue.

OpenStudy (abb0t):

No, use it to find your \[x_i \] remember \[x_i = a + (∆x)i \]

OpenStudy (anonymous):

Ok so that becomed 0+i(2/n) = 2i/n right?

OpenStudy (abb0t):

Yep. There are three or 4 formulas u can use. I suggest you memorize them if you're not given them on an exam. i will draw them out for you

OpenStudy (anonymous):

And then do u do u f(xi) = x^2+5 = (2i/n)^2 + 5 = (4i^2/n^2) + 5

OpenStudy (abb0t):